Let {x1,x2,..., Xn} be a ..., Xn} be a sample from Bernoulli(p). Find an unbiased estimator for p?.Let {x1, x2, ..., xn} be a sample from Bernoulli(p). Find an unbiased estimator for p^2 .
Answer #1

Qus: find the unbaised estimator of p^2 for Bernoulli distribution

Solution:

Solution - be a sample from Rerolls (o) We have to find an unlaised for x=0, estizator for p² We know that if X ~ Berbullillaagain we know that Veriance of Randon varitte & which fallow Berboulli (f) is equal to so (where gitt Vas (Y) = P(1-P) We kno(7) - p?+ PC1-2) 12 р? + Р pe 1) 1 ܠܐ E (X² - Do = (1-1) p² + b Elx²) p. (1-4) p² (1-1) p² (p= E(8) b Ebx) = GELY) E(72) - E(